# Topological Data Analysis for Neural Network Analysis: A Comprehensive Survey

### 10 Dec 2023 v1

@@ 1 Introduction
Over the past few years, deep learning has consolidated its position as the most successful branch of artificial intelligence. With the continuous growth in computational capacity, neural networks have expanded in size and complexity, enabling them to effectively tackle increasingly challenging problems. However, their increased capacity has made it more challenging to comprehend essential properties of the networks such as their interpretability, generalization ability, or suitability for specific problems. From both theoretical and practical standpoints, this is undesirable, especially in critical contexts where AI decisions could lead to catastrophic consequences, such as medical diagnosis (Yang et al., 2021) or autonomous driving (Wäschle et al., 2022), among others.
Topological Data Analysis (TDA) has emerged as a subfield of algebraic topology that offers a framework for gaining insights into the shape of data in a broad sense. Topological data analysis has found application across a wide array of experimental science disciplines, spanning from biomedicine (Skaf and Laubenbacher, 2022) to finance (Gidea and Katz, 2018), among numerous others. One of its most prolific areas of application is machine learning, particularly in the domain of deep learning. A basic introduction to topological machine learning can be found in Hensel et al. (2021). Topological data analysis, specifically homology, persistent homology and Mapper, has been used to analyze various aspects of neural networks. Broadly, these aspects can be categorized in the following four groups: 1. Structure of the neural network; 2. Input and output spaces; 3. Internal representations and activations; 4. Training dynamics and loss functions.
@@
Figure 1: Diagram showing the usual lifecycle of a neural network $\mathcal{N}$. First, an architecture $a(\mathcal{N})$ is selected based on the task to be solved. This architecture is independent of the learned parameters $\theta(\mathcal{N})$ or the specific input data used to train or test the network, denoted $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{test}}$, respectively. Second, the architecture is trained (T) using a specific training algorithm $\mathcal{A}$, which generally minimizes the empirical risk of a loss function $\mathcal{L}$ evaluated in the training dataset $\mathcal{D}_{\text{train}}$. Once the network is trained, inference (I) is performed with data coming from the same distribution $\mathbb{P}$ from which the training data were sampled. For trained neural networks, input and output spaces gather several interesting structures, such as decision regions and boundaries for classification problems or latent spaces for generative models, among others. Each dashed box is related to one of the categories, labeled $1$ to $4$, in which topological data analysis has been used to analyze neural networks. The categories are the following: (1) structure of the neural network; (2) input and output spaces; (3) internal representations and activations; (4) training dynamics and loss functions. The leftmost box contains the training part of the lifecycle of a neural network and is related to category 4. The central box contains the neural network and is concerned with categories 1 and 3. The rightmost box contains the decision regions and boundaries of the output space of a neural network after training, which are related to category 2.
@@ 2.1 Notation
The support of a function $f\colon X\to\mathbb{R}$, denoted $\text{supp}(f)$, is the set of elements of $X$ that are not sent to zero, or most commonly the closure of this set if $X$ is a topological space. Similarly, for a probability distribution $\mathbb{P}$ associated with a random variable $X$, the support $\text{supp}(\mathbb{P})$ is the smallest closed set $S$ of real numbers such that $P(X\in S)=1$. We also call support of $\mathbb{P}$ the set of elements of the sample space mapped to $\text{supp}(\mathbb{P})$. The letter $\mathbb{E}$ is used to denote expectation of a random variable.
@@ 2.2 Deep Learning
One of the main difficulties in this approach is that the data distribution $\mathbb{P}_{(X,Y)}$ is often unknown, and we only have access to a finite sample $\mathcal{D}=\{\left(x_{i},y_{i}\right)\}_{i=1}^{m}$ from it, that we assume independent and identically distributed (i.i.d.) with distribution $\mathbb{P}_{(X,Y)}^{m}$. Therefore, instead of trying to minimize directly the risk $\mathcal{R}(f)$, one tries to minimize the empirical risk ^$\hat{\mathcal{R}}_{\mathcal{D}_{\text{train}}}(f)$ for a subset $\mathcal{D}_{\text{train}}\subseteq\mathcal{D}$, called training dataset, given by
^$\hat{\mathcal{R}}_{\mathcal{D}_{\text{train}}}(f)=\frac{1}{m}\sum_{i=1}^{m}\mathcal{L}\left(f,x_{i},y_{i}\right),$ (1)
@@
We say that a prediction task is a regression task if $\mathcal{X}$ is a Euclidean space $\mathbb{R}^{d}$ for some $d\in\mathbb{N}_{>0}$ and $\mathcal{Y}=\mathbb{R}$. On the other hand, we say that a prediction task is a classification task if $\mathcal{X}$ is still a Euclidean space $\mathbb{R}^{d}$ but $\mathcal{Y}$ is a finite set. Without loss of generality, we assume that $\mathcal{Y}=[k]$ with $k\in\mathbb{N}_{\geq 2}$, and that $\text{supp}\left(\mathbb{P}_{Y}\right)=\mathcal{Y}$. If $k=2$, then we call it a binary classification task. For a broader introduction to machine learning tasks, we refer the reader to Goodfellow et al. (2016, Section 5.1.1).
Generative tasks are not as straightforward to define as prediction tasks, and formal definitions can vary depending on the problem or the solving method. In generative tasks, the goal is to generate synthetic data that are indistinguishable from some source data. We assume that source data live in an ambient space $\mathcal{X}$ and follow a distribution $\mathbb{P}_{X}$. The objective is to learn a function $G\colon\mathcal{Z}\to\mathcal{X}$, in our case given by a neural network $\mathcal{G}$, called generator, from a probability space $\mathcal{Z}$ with distribution $\mathbb{P}_{Z}$ to $\mathcal{X}$, such that the probability distribution $G\left(\mathbb{P}_{Z}\right)$ on $\mathcal{X}$ induced by $G$ is as close as possible to $\mathbb{P}_{X}$. To measure the quality of a generator function, one uses a loss function, that in its simplest form is a differentiable function $\mathcal{L}\colon\mathcal{M}\left(\mathcal{Z},\mathcal{X}\right)\times\mathcal{X}\to\mathbb{R}$. We follow the same methodology as above to obtain functions $G$, by defining training algorithms $\mathcal{A}$ that minimize, given a dataset $\mathcal{D}_{\text{train}}$ with distribution $\mathbb{P}_{X}^{m}$, the empirical risk ^$\hat{\mathcal{R}}_{\mathcal{D}_{\text{train}}}(\theta)$ for a hypothesis set $\mathcal{F}$ consisting of neural networks.
In this survey we mainly consider fully connected feedforward neural networks (FCFNN). Fully connected feedforward neural networks follow a basic structure in which most of the current, more complex, models, can be expressed. A FCFNN $\mathcal{N}$ is a parameterized function that can be specified by an architecture $a(\mathcal{N})=(N,\varphi)$, where $N=(N_{0},\ldots,N_{L})\in\mathbb{N}^{L+1}$ and $\varphi=\left(\varphi^{(l)}\right)_{l=1}^{L}$ is a tuple of differentiable functions $\varphi^{(l)}\colon\mathbb{R}\to\mathbb{R}$, where differentiability is meant except perhaps on a set of zero measure, and a set of parameters $\theta(\mathcal{N})=\left(\theta^{(l)}\right)_{l=1}^{L}$. Regarding the architecture, $L$ is the number of layers of the neural network, while $N_{0}$, $N_{L}$, and $N_{l}$ for $l\in[L-1]$ are the number of neurons of the input, output, and $l$-th hidden layers, respectively, and $\varphi^{(l)}$ is the activation function of the layer $l$. Regarding the parameters, $\theta^{(l)}=\left(W^{(l)},b^{(l)}\right)$ are the parameters of the layer $l$, belonging to $\mathbb{R}^{N_{l}\times N_{l-1}}\times\mathbb{R}^{N_{l}}$ for $l\in[L]$, and known as weights and biases, respectively.
@@
where each component $x_{i}$ of $x$ is associated with the vertex $v_{0}^{i}$, and each intermediate function value $\phi^{(l)}_{\mathcal{N}}(x)_{i}$ is associated with the vertex $v_{l}^{i}$. Given an input value $x$, we say that the values $\phi^{(l)}_{\mathcal{N}}(x)_{i}$ are the activations of the vertices $v_{l}^{i}$. A visual example of a FCFNN is shown in Figure 2.
@@
Given a specific tuple $a=\left(N,\varphi\right)$ representing an architecture for a FCFNN and a possible set of parameters $\Theta\subseteq\prod_{l\in[L]}\mathbb{R}^{\mathbb{N}_{l}\times\mathbb{N}_{l-1}}\times\mathbb{R}^{N_{l}}$, we denote the hypothesis set of all the neural networks represented by the architecture $a$ with parameters in $\Theta$ as
@@
Note that, fixing a dataset $\mathcal{D}$, an architecture $a$, and a set of possible parameters $\Theta$, the composition of the empirical risk with the neural network functions in $\mathcal{F}_{a,\Theta}$ can be seen as a function ^$\hat{\mathcal{R}}_{\mathcal{D}_{\text{train}}}(\theta)$ of the parameters $\theta$. Usually, training algorithms for neural networks, given an initial set of parameters $\theta^{(0)}\in\Theta$, generate a discrete weight trajectory $\left(\theta^{(i)}\right)_{i=1}^{N}$ by minimizing the empirical risk ^$\hat{\mathcal{R}}_{\mathcal{D}_{\text{train}}}(\theta)$ iteratively with respect to the parameters $\theta$ until some stopping criteria are satisfied. The output of the algorithm is one of the weights $\theta^{(i)}$ generated in the trajectory, usually the last one, $\theta^{(N)}$. In this survey, we assume that neural networks are trained in this way.
@@
A special type of FCFNN is a convolutional neural network (CNN), which is a FCFNN with a specific architecture designed to work with data that have a grid-like structure, especially images. A CNN is one of the most common type of neural networks used in computer vision. CNNs implement two functions between layers that impose restrictions on the set of weights and on the activation functions: convolutional layers and pooling operators. A convolutional layer is a layer whose outputs are the result of performing convolutions between the values associated to the neurons of its input layer rearranged as a grid and a filter tensor that imposes a set of equalities on the coefficients of the weight matrices. In one of most typical scenarios, a convolutionSal layer between layer $i-1$ and $i$ is a transformation between the activations of $V_{i-1}$ and $V_{i}$ in such a way $V_{i-1}$ and $V_{i}$ are ordered in grid structures of size $h^{(i-1)\times}w^{(i-1)}\times c^{(i-1)}$ and $h^{(i)\times}w^{(i)}\times c^{(i)}$, respectively. We say that a layer $i$ arranged in this grid structure has height $h^{(i)}$, width $w^{(i)}$, and $c^{i}$ channels, where each channel is defined as the subgrid obtained by fixing one value in $\left[c^{i}\right]$ for the third dimension. The convolutional layer transformation is produced by (discretely) convolving $c^{(i)}$ filter tensors $C^{(i)}_{j}$, $j\in\left[d^{(i)}\right]$ of size $h\times w\times c^{(i-1)}$, where $h\times w$ is called convolution size, with the activations in the grid structure of $V_{i-1}$, obtaining $c^{(i)}$ convolved grids of size $h^{(i)}\times w^{(i)}$, that are then concatenated to obtain the activations of $V_{i}$. Pooling operators are similar operations that allow to reduce the dimensionality of its input layer. Note that these operations reduce or mantain the size of the input, and that not necessarily all values in the input affect each output of the layer. Given a neuron of the layer, the region of the input that affects its activations is called its receptive field. For a thorough introduction to CNNs, we refer the reader to Goodfellow et al. (2016, Chapter 9).
@@
Fully connected feedforward networks define functions $\phi_{\mathcal{N}}\colon\mathbb{R}^{N_{0}}\to\mathbb{R}^{N_{L}}$, which is optimal whenever $\mathcal{X}$ and $\mathcal{Y}$ are subsets of Euclidean spaces. However, there are some specific tasks, mostly classification ones, in which we deal with a finite set of points in the output space, that is, there exists $k\in\mathbb{N}$ such that $\mathcal{Y}=[k]$. In these cases, we select a projection function $\pi\colon\mathbb{R}^{N_{L}}\to[k]$ that projects the outputs of the neural network $\mathcal{N}$ into the finite set of points $[k]$ and we use $\pi\circ\phi_{\mathcal{N}}\colon\mathbb{R}^{N_{0}}\to[k]$ as a model. In this context, the decision region of a label $y\in[k]$ is the set of inputs $x$ such that $\pi(\phi_{\mathcal{N}}(x))=y$. To train the neural network, it is common to choose surrogate loss functions $\mathcal{L}^{\text{surr}}$ that depend directly on the output of $\phi_{\mathcal{N}}$. In classification tasks, the most common configuration for a problem with $\mathcal{X}$ having $\mathbb{R}^{d}$ as an ambient space and $\mathcal{Y}=[k]$ is to choose neural networks with $N_{0}=d$, $N_{L}=k$, projection $\pi(x_{1},\ldots,x_{k})=\text{argmax}_{i\in[k]}\,x_{i}$, and surrogate loss function $\mathcal{L}^{\text{surr}}$ depending only on $\phi_{\mathcal{N}}$ like the categorical cross entropy loss CE, that, in its most basic form, looks like
@@
For neural networks with the configuration $N_{L}=k$ and $\pi(x_{1},\ldots,x_{k})=\text{argmax}_{i\in[k]}\,x_{i}$, there is an ambiguity whenever the argmax is not unique, i.e., when there are at least two indices $i,j\in[k]$ such that $x_{i}=x_{j}$. In this case, the projection function must define a deterministic way to choose one of the equal-valued indices. Given a neural network $\mathcal{N}$ as the previous one, the points $x\in\mathcal{X}$ for which there exists an ambiguity is called decision boundary of $\mathcal{N}$. Formally,
$\mathcal{B}\left(\mathcal{N}\right)=\left\{x\in\text{Dom}(\phi_{\mathcal{N}}):\exists i,j\in[N_{L}]\text{ such that }\phi_{\mathcal{N}}(x)_{i}=\phi_{\mathcal{N}}(x)_{j}=\max_{\iota\in[k]}\phi_{\mathcal{N}}(x)_{\iota}\right\}.$ (4)
Sometimes, decision regions are studied without the decision boundary, that is, $(\pi\circ\phi_{\mathcal{N}})^{-1}(y)\smallsetminus\mathcal{B}(\mathcal{N})$ is used. Figure 3 shows examples of decision boundaries and regions for classification problems with three labels.
@@
Figure 3: Decision regions and boundaries for three different classification problems with three labels. The black lines represent decision boundaries given by FCFNNs. The different decision regions are separated by decision boundaries. Decision regions and boundaries give valuable information on the neural network used in each case. Classification problems are ordered from left to right by increasing complexity. For the left and central classification problems, the decision regions and boundaries are simple, in the sense that they seem to properly classify the inputs of the domain without visible outliers or strange regions. However, the neural network for the central classification problem seems to have a more complex output, since the blue decision region has one hole that does not exist in the first case. This could be, for example, an indicator of the inherent difficulty of the problem. The right decision regions and boundary are more complicated. The blue decision region has two connected components and one hole, and the red and green regions have two odd protuberances which appeared due to outliers in the training data. Although protuberances cannot be detected by the use of usual topological techniques, the extra connected component is detected by the homology groups of the blue decision region. See Section 2.3 for more details.

### 03 Jan 2024 v2

@@ 1 Introduction
Over the past few years, deep learning has consolidated its position as the most successful branch of artificial intelligence. With the continuous growth in computational capacity, neural networks have expanded in size and complexity, enabling them to effectively tackle progressively difficult problems. However, their increased capacity has made it more challenging to comprehend essential properties of the networks such as their interpretability, generalization ability, or suitability for specific problems. From both theoretical and practical standpoints, this is undesirable, especially in critical contexts where AI decisions could lead to catastrophic consequences, such as medical diagnosis (Yang et al., 2021) or autonomous driving (Wäschle et al., 2022), among others.
Topological Data Analysis (TDA) has emerged as a subfield of algebraic topology that offers a framework for gaining insights into the shape of data in a broad sense. Topological data analysis has found application across a wide array of experimental science disciplines, spanning from biomedicine (Skaf and Laubenbacher, 2022) to finance (Gidea and Katz, 2018), among numerous others. One of its most prolific areas of application is machine learning, particularly in the domain of deep learning. A basic introduction to topological machine learning can be found in Hensel et al. (2021). Topological data analysis, specifically homology, persistent homology and Mapper, has been used to analyze various aspects of neural networks. Broadly, these aspects can be categorized in the following four groups: 1. Structure of neural networks; 2. Input and output spaces; 3. Internal representations and activations; 4. Training dynamics and loss functions.
@@
Figure 1: Diagram showing the usual lifecycle of a neural network $\mathcal{N}$. First, an architecture $a(\mathcal{N})$ is selected based on the task to be solved. This architecture is independent of the learned parameters $\theta(\mathcal{N})$ or the specific input data used to train or test the network, denoted $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{test}}$, respectively. Second, the architecture is trained (T) using a specific training algorithm $\mathcal{A}$, which generally minimizes the empirical risk of a loss function $\mathcal{L}$ evaluated on the training dataset $\mathcal{D}_{\text{train}}$. Once the network is trained, inference (I) is performed with data coming from the same distribution $\mathbb{P}$ from which the training data were sampled. For trained neural networks, input and output spaces gather several interesting structures, such as decision regions and boundaries for classification problems or latent spaces for generative models, among others. Each dashed box is related to one of the categories, labeled $1$ to $4$, in which topological data analysis has been used to analyze neural networks. The categories are the following: (1) Structure of the neural network; (2) Input and output spaces; (3) Internal representations and activations; (4) Training dynamics and loss functions. The leftmost box contains the training part of the lifecycle of a neural network and is related to category 4. The central box contains the neural network and is concerned with categories 1 and 3. The rightmost box contains the decision regions and boundaries of the output space of a neural network after training, which are related to category 2.
@@ 2.1 Notation
The support of a function $f\colon X\to\mathbb{R}$, denoted $\text{supp}(f)$, is the set of elements of $X$ that are not sent to zero, or most commonly the closure of this set if $X$ is a topological space. Similarly, for a probability distribution $\mathbb{P}$ associated with a random variable $X$, the support $\text{supp}(\mathbb{P})$ is the smallest closed set $S$ of real numbers such that $\mathbb{P}(X\in S)=1$. We also call support of $\mathbb{P}$ the set of elements of the sample space mapped to $\text{supp}(\mathbb{P})$ by $X$. The letter $\mathbb{E}$ is used to denote expectation of a random variable.
@@ 2.2 Deep learning
One of the main difficulties in this approach is that the data distribution $\mathbb{P}_{(X,Y)}$ is often unknown, and we only have access to a finite sample $\mathcal{D}=\{\left(x_{i},y_{i}\right)\}_{i=1}^{m}$ from it, that we assume independent and identically distributed (i.i.d.) with distribution $\mathbb{P}_{(X,Y)}^{m}$. Therefore, instead of trying to minimize directly the risk $\mathcal{R}(f)$, one tries to minimize the empirical risk ^$\widehat{\mathcal{R}}_{\mathcal{D}_{\text{train}}}(f)$ for a subset $\mathcal{D}_{\text{train}}\subseteq\mathcal{D}$, called training dataset, given by
^$\widehat{\mathcal{R}}_{\mathcal{D}_{\text{train}}}(f)=\frac{1}{m}\sum_{i=1}^{m}\mathcal{L}\left(f,x_{i},y_{i}\right),$ (1)
@@
We say that a prediction task is a regression task if $\mathcal{X}$ is a Euclidean space $\mathbb{R}^{d}$ for some $d\in\mathbb{N}_{>0}$ and $\mathcal{Y}=\mathbb{R}$, and we say that a prediction task is a classification task if $\mathcal{X}$ is a Euclidean space $\mathbb{R}^{d}$ and $\mathcal{Y}$ is a finite set. In the latter case we assume, without loss of generality, that $\mathcal{Y}=[k]$ with $k\in\mathbb{N}_{\geq 2}$ and that $\text{supp}\left(\mathbb{P}_{Y}\right)=\mathcal{Y}$. If $k=2$, then we call it a binary classification task. For a broader introduction to machine learning tasks, we refer the reader to Goodfellow et al. (2016, Section 5.1.1).
Generative tasks are not as straightforward to define as prediction tasks, and formal definitions can vary depending on the problem or the solving method. In generative tasks, the purpose is to generate synthetic data that are indistinguishable from source data. We assume that source data live in an ambient space $\mathcal{X}$ and follow a distribution $\mathbb{P}_{X}$. Some models explicitly approximate the given probability distribution, allowing to directly sample from it, while other models learn a mechanism to generate new examples without explicitly describing the data distribution. We refer to Prince (2023, Section 1.2.1) for a detailed explanation of generative models.
In this survey we mainly consider fully connected feedforward neural networks (FCFNN). FCFNNs follow a basic structure in which most of the current, more complex, models, can be expressed. A FCFNN is a parameterized function $\mathcal{N}$ consisting of:
  • (i)
    An architecture $a(\mathcal{N})=(N,\varphi)$, where $N=(N_{0},\ldots,N_{L})\in\mathbb{N}^{L+1}$ and $\varphi=\left(\varphi^{(l)}\right)_{l=1}^{L}$ is a tuple of differentiable functions $\varphi^{(l)}\colon\mathbb{R}\to\mathbb{R}$. Here differentiability is meant except perhaps on a set of zero measure. The number $L$ counts the layers of the network, while $N_{0}$, $N_{L}$, and $N_{l}$ for $l\in[L-1]$ are the numbers of neurons of the input, output, and $l$-th hidden layers, respectively, and $\varphi^{(l)}$ is the activation function of layer $l$.
  • (ii)
    A set of parameters $\theta(\mathcal{N})=\left(\theta^{(l)}\right)_{l=1}^{L}$, where $\theta^{(l)}=\left(W^{(l)},b^{(l)}\right)$ are the parameters of layer $l$, belonging to $\mathbb{R}^{N_{l}\times N_{l-1}}\times\mathbb{R}^{N_{l}}$ for $l\in[L]$ and known as weights and biases, respectively.
@@
where each component $x_{i}$ of $x$ is associated with the vertex $v_{0}^{i}$, and each intermediate function value $\phi^{(l)}_{\mathcal{N}}(x)_{i}$ is associated with the vertex $v_{l}^{i}$. Given an input value $x$, we say that the values $\phi^{(l)}_{\mathcal{N}}(x)_{i}$ are activations of the vertices $v_{l}^{i}$. A visual example of a FCFNN is shown in Figure 2.
@@
Given a specific tuple $a=\left(N,\varphi\right)$ representing an architecture for a FCFNN and a subset $\Theta\subseteq\prod_{l\in[L]}\mathbb{R}^{\mathbb{N}_{l}\times\mathbb{N}_{l-1}}\times\mathbb{R}^{N_{l}}$, we denote the hypothesis set of all the neural networks represented by the architecture $a$ with parameters in $\Theta$ as
@@
Note that, fixing a dataset $\mathcal{D}$, an architecture $a$, and a set of possible parameters $\Theta$, the composition of the empirical risk with the neural network functions in $\mathcal{F}_{a,\Theta}$ can be seen as a function ^$\widehat{\mathcal{R}}_{\mathcal{D}_{\text{train}}}(\theta)$ of the parameters $\theta$. Usually, training algorithms for neural networks, given an initial set of parameters $\theta^{(0)}\in\Theta$, generate a discrete weight trajectory $\left(\theta^{(i)}\right)_{i=1}^{N}$ by minimizing the empirical risk ^$\widehat{\mathcal{R}}_{\mathcal{D}_{\text{train}}}(\theta)$ iteratively with respect to the parameters $\theta$ until some stopping criteria are satisfied. The output of the algorithm is one of the weights $\theta^{(i)}$ generated in the trajectory, usually the last one, $\theta^{(N)}$. In this survey, we assume that neural networks are trained in this way.
@@
A special type of FCFNN is a convolutional neural network (CNN), which is a FCFNN with a specific architecture designed to work with data that have a grid-like structure, especially images. A CNN is one of the most common type of neural networks used in computer vision. CNNs implement two functions between layers that impose restrictions on the set of weights and on the activation functions: convolutional layers and pooling operators. A convolutional layer is a layer whose outputs are the result of performing convolutions between the values associated to the neurons of its input layer rearranged as a grid and a filter tensor that imposes a set of equalities on the coefficients of the weight matrices. In one of the most typical scenarios, a convolutional layer between layer $i-1$ and $i$ is a transformation between the activations of $V_{i-1}$ and $V_{i}$ in such a way $V_{i-1}$ and $V_{i}$ are ordered in grid structures of size $h^{(i-1)\times}w^{(i-1)}\times c^{(i-1)}$ and $h^{(i)\times}w^{(i)}\times c^{(i)}$, respectively. We say that a layer $i$ arranged in this grid structure has height $h^{(i)}$, width $w^{(i)}$, and $c^{i}$ channels, where each channel is defined as the subgrid obtained by fixing one value in $\left[c^{i}\right]$ for the third dimension. The convolutional layer transformation is produced by (discretely) convolving $c^{(i)}$ filter tensors $C^{(i)}_{j}$, $j\in\left[d^{(i)}\right]$ of size $h\times w\times c^{(i-1)}$, where $h\times w$ is called convolution size, with the activations in the grid structure of $V_{i-1}$, obtaining $c^{(i)}$ convolved grids of size $h^{(i)}\times w^{(i)}$, that are then concatenated to obtain the activations of $V_{i}$. Pooling operators are similar operations aiming to reduce the dimensionality of the input layer, which reduce or mantain the size of the input, and not necessarily all values in the input affect each output of the layer. Given a neuron, the region of the input that affects its activations is called its receptive field. For a thorough introduction to CNNs, we refer the reader to Goodfellow et al. (2016, Chapter 9).
@@
Fully connected feedforward networks define functions $\phi_{\mathcal{N}}\colon\mathbb{R}^{N_{0}}\to\mathbb{R}^{N_{L}}$, which is optimal whenever $\mathcal{X}$ and $\mathcal{Y}$ are subsets of Euclidean spaces. However, there are specific tasks, mostly classification ones, in which one deals with a finite set of points in the output space, that is, there exists $k\in\mathbb{N}$ such that $\mathcal{Y}=[k]$. In these cases, one selects a projection function $\pi\colon\mathbb{R}^{N_{L}}\to[k]$ that projects the outputs of the neural network $\mathcal{N}$ into the finite set of points $[k]$ and one uses $\pi\circ\phi_{\mathcal{N}}\colon\mathbb{R}^{N_{0}}\to[k]$ as a model. In this context, the decision region of a label $y\in[k]$ is the set of inputs $x$ such that $\pi(\phi_{\mathcal{N}}(x))=y$. To train the neural network, it is common to choose surrogate loss functions $\mathcal{L}^{\text{surr}}$ that depend directly on the output of $\phi_{\mathcal{N}}$. In classification tasks, the most common configuration for a problem with $\mathcal{X}$ having $\mathbb{R}^{d}$ as an ambient space and $\mathcal{Y}=[k]$ is to choose neural networks with $N_{0}=d$, $N_{L}=k$, projection $\pi(x_{1},\ldots,x_{k})=\text{argmax}_{i\in[k]}\,x_{i}$, and surrogate loss function $\mathcal{L}^{\text{surr}}$ depending only on $\phi_{\mathcal{N}}$ like the categorical cross entropy loss CE, that, in its most basic form, looks like
@@
For neural networks with the configuration $N_{L}=k$ and $\pi(x_{1},\ldots,x_{k})=\text{argmax}_{i\in[k]}\,x_{i}$, there is an ambiguity when the argmax is not unique, i.e., when there are at least two indices $i,j\in[k]$ such that $x_{i}=x_{j}$. In this case, the projection function must define a deterministic way to choose one of the equal-valued indices. Given a neural network $\mathcal{N}$ as the previous one, the points $x\in\mathcal{X}$ for which there exists an ambiguity is called decision boundary of $\mathcal{N}$. Formally,
$\mathcal{B}\left(\mathcal{N}\right)=\Big{\{}x\in\text{Dom}(\phi_{\mathcal{N}}):\exists i,j\in[N_{L}]\text{ such that }\phi_{\mathcal{N}}(x)_{i}=\phi_{\mathcal{N}}(x)_{j}=\max_{\iota\in[k]}\phi_{\mathcal{N}}(x)_{\iota}\Big{\}}.$ (4)
Sometimes, decision regions are studied without the decision boundary; in other words, $(\pi\circ\phi_{\mathcal{N}})^{-1}(y)\smallsetminus\mathcal{B}(\mathcal{N})$ is used. Figure 3 shows examples of decision boundaries and regions for classification problems with three labels.
@@
Figure 3: Decision regions and boundaries for three different classification problems with three labels. The black lines represent decision boundaries given by FCFNNs. The different decision regions are separated by decision boundaries. Decision regions and boundaries give valuable information on the neural network used in each case. Classification problems are ordered from left to right by increasing complexity. For the left and central classification problems, the decision regions and boundaries are simple, in the sense that they seem to properly classify the inputs of the domain without visible outliers or strange regions. However, the neural network for the central classification problem seems to have a more complex output, since the blue decision region has one hole that does not exist in the first case. This could be, for example, an indicator of the inherent difficulty of the problem. The right decision regions and boundary are more complicated. The blue decision region has two connected components and one hole, and the red and green regions have two odd protuberances which appeared due to outliers in the training data. Although protuberances cannot be detected by the use of usual topological techniques, the extra connected component is detected by the homology of the blue decision region. See Section 2.3 for more details.